# 1]FAIR at Meta

### v1

@@ 1 Introduction
In psychology, the dual process theory Wason and Evans (1974) explains that thought can emerge through two distinct processes. Typically, one process is implicit (automatic and unconscious) and the other process is explicit (controlled and conscious). The famous book Thinking, Fast and Slow Kahneman (2017) explores the concept of dual process theory extensively. It describes two different systems of thinking. System 1, corresponding to the implicit process, is fast, intuitive, and emotional. System 2, corresponding to the explicit process, is slower, more deliberative, and more logical. Kahneman applies this theory to a wide range of cognitive and behavioral activities to illustrate how these two systems influence our decision making and reasoning. As outlined in the book, System 1 can lead to biases and is prone to systematic errors due to the lack of voluntary control. As a counterbalance, System 2 is more suitable for tasks that require careful analysis, consideration, and planning.
Transformers Vaswani et al. (2017), the sequence modeling tool that serves as the cornerstone of foundation models in various domains including Large Language models (LLMs) Dosovitskiy (2020); Baevski et al. (2020); Radford et al. (2021); Touvron et al. (2021); Hsu et al. (2021); Touvron et al. (2023); Dubey et al. (2024), have been widely used in many works to approach reasoning and planning problems, see e.g.,  Zhou et al. (2022); Kojima et al. (2022); Pallagani et al. (2022); Valmeekam et al. (2023a); Chen et al. (2024); Gundawar et al. (2024); Wang and Zhou (2024). Interestingly, the dual process theory generalizes to artificial intelligence, where we can categorize the reasoning modes of Transformer into fast and slow. In fast mode, a Transformer will directly output a final solution without any reasoning steps, whereas the intermediate steps of thinking, such as a search trace for finding a short path, will be generated along with the plan in slow mode. The two inference modes share similar pros and cons with the two thinking systems inherent in us. As discussed in previous works Wei et al. (2022b); Valmeekam et al. (2023b); Lehnert et al. (2024); Gandhi et al. (2024); Saha et al. (2024), models operating in fast mode have a lower computational cost and allow for a quicker response, yet they fall short in accuracy and optimality compared with slow mode models. This raises the question:
@@
We first apply our framework to train an encoder-decoder Transformer model to solve pathfinding problems, where the trace is generated by the A* search algorithm. We consider two domains: the Maze navigation and the Sokoban game as in Lehnert et al. (2024), where we use the same tokenization scheme. Interestingly, we have found that these problems are challenging for state-of-the-art LLMs like o1-preview and o1-mini, where the output path often breaks into the walls (See Section 13 for an example). In each reasoning mode, Dualformer outperforms the established baselines, achieving stronger results in both the solved rate and optimal rate. Moreover, Dualformer significantly enhances the diversity of the generated plans by identifying a greater variety of unique paths that reach the goal. Notably, Dualformer is also efficient even when working in slow mode, generating much shorter reasoning traces than the baseline model. Next, we apply our framework to finetune LLMs for answering math questions. Following Yu et al. (2023), the training examples are taken from the MATH dataset Hendrycks et al. (2021) where answers are rewritten by a Llama-3.1-70B-Instruct model to include detailed intermediate steps. Likewise, the obtained LLMs demonstrate enhanced efficacy and efficiency.
@@ 2 Related Work
Cognitive Science & Psychology
Wason and Evans (1974) first introduces the dual-process theory, which is then extensively explored and explained in Kahneman (2017). This framework has underpinned extensive research in decision-making across various fields, including psychology and behavioral economics Thaler and Sunstein (2008); Povey et al. (2000); Tversky and Kahneman (1974); Kahneman (2017). In cognitive science, the distinction between System 1 and System 2 has facilitated a deeper understanding of how humans process information and make decisions. Research has shown that System 1 operates automatically and quickly, with little or no effort and no sense of voluntary control, often relying on heuristics that lead to biases Tversky and Kahneman (1974). Conversely, System 2 allocates attention to effortful mental activities and is associated with the subjective experience of agency, choice, and concentration.
@@
Tremendous efforts have been made to enhance the capability of Transformer-based models to plan and reason over a long horizon. Two main types of approaches have been developed. The first type leverages existing LLMs. For instance, researchers taught LLMs to call external existing symbolic solvers, such as those found in Ahn et al. (2022); Besta et al. (2024); Sel et al. (2023); He-Yueya et al. (2023); Liu et al. (2023); Silver et al. (2024). Pallagani et al. (2022, 2024) investigate the finetuning of LLMs to use a symbolic solver, Schick et al. ; Hao et al. (2024) fine-tune LLMs to use external tools, and Hao et al. propose to finetune LLMs to do reasoning and planning within a world model. Among them, several works try to integrate System-1 and System-2 thinking into LLM reasoning. Weston and Sukhbaatar (2023) proposed the System 2 Attention (S2A), a more deliberate attention mechanism aimed at improving the reasoning capabilities of large language models by reducing their susceptibility to spurious correlations in the context. Yu et al. (2024) distill System 2 output into a System 1 model, which aims to compile higher-quality outputs from System 2 techniques back into LLM generations without intermediate reasoning token sequences.
The second type aims to train Transformers from scratch to plan and reason independently Lehnert et al. (2024); Saha et al. (2024); Gandhi et al. (2024), using task-specific language. Both Lehnert et al. (2024) and Gandhi et al. (2024) teach LLMs to search by representing the search process in language, Lehnert et al. (2024) developed Searchformer that is trained to mimic the A* search process for pathfinding problems. Gandhi et al. (2024) applied their model to the Countdown game. In the concurrent work Saha et al. (2024), the authors trained two separate models and manage them by an external meta-controller. There, one model generates rapid traceless response, while another one responds slower but with trace. Our research is closely related to the above work, but has some key differences. Unlike Lehnert et al. (2024); Gandhi et al. (2024) which do not modify the reasoning trace in the training data, our approach involves randomizing the reasoning trace. Unlike Saha et al. (2024), we do not use any explicit controller nor use two networks for each mode. Instead, we integrate both the fast and the slow mode functionalities into a single model.
@@
Large language models (LLMs) have been utilized for synthetic data generation in various domains. For instance, Wei et al. (2022a); Longpre et al. (2023); Taori et al. (2023) introduced a synthetic instruction dataset by sampling from diverse templates containing natural language instructions that outline specific tasks. This approach has also been applied to the visual domain Liu et al. (2024b, a); Zhu et al. (2023); Brooks et al. (2023); Peng et al. (2023). To improve the performance of LLM to answer math questions, Yu et al. (2023) developed a method to rewrite, verify, and augment the original MATH dataset Hendrycks et al. (2021) using specialized prompts. Similar methodologies have been further explored in other studies, including those by  Yuan et al. (2023); Luo et al. (2023); Lee et al. (2023); Yue et al. (2023).
@@ 4 Structured Trace Dropping and Randomized Training
Searchformer proposed by Lehnert et al. (2024) has proven efficacy in addressing a variety of complex decision-making tasks. However, it still suffers from two important limitations. Firstly, the model only operates in slow mode and outputs lengthy reasoning chains, which significantly increases the inference time. While this can be reduced by bootstrapping Lehnert et al. (2024), an iterative refining technique that consists of cycles of rollouts followed by fine-tuning, such a procedure incurs significant extra demand on computational resources. Secondly, Searchformer struggles to generate diverse solutions, where identical rollouts are frequently sampled [^1] We note that the lack of diversity is not due to the limitation of training data. For all our experiments, the search traces in the training datasets are generated by non-deterministic A* algorithm that randomly breaks cost ties.. For example, across 1000 $30\times 30$ maze problems we have tested, Searchformer’s reasoning chain contains more than 1500 tokens on average, and can only find $7.6$ unique feasible paths out of $64$ responses (see Section 5.1.2).
@@ 5.1 Navigation Tasks: Maze & Sokoban
We evaluate whether a model generates correct and optimal plans using two metrics: 1-Solved-64 and 1-Optimal-64. Namely, for each evaluation task (e.g. one maze or one Sokoban game), we randomly sample 64 responses from a trained model. Each response is parsed and evaluated regardless of the generated trace part. If any of the 64 plans is correct, i.e. is feasible and reaches the goal location, this task is labelled as success for the 1-Solved-64 metric. If any of the 64 plans is optimal, this task is labelled as success for the 1-Optimal-64 metric. We repeat such process for 1000 unseen evaluation tasks and report the average success rate. To investigate the robustness of each method, we also report metrics 3-Solved-64 and 3-Optimal-64 where a task is labelled as success if at least 3 plans are correct or optimal. Additionally, we consider the Success Weighted by Cost (Wu et al., 2019, SWC) that measures the quality of the resulting plans in terms of their costs, aggregated over individual responses. More precisely: $\text{SWC}=\tfrac{1}{64}\sum_{i=1}^{64}\mathbb{I}(\text{plan}\,i\,\text{is correct})\cdot\tfrac{c^{*}}{c_{i}}$, where $\mathbb{I}$ is the indicator function, $c^{*}$ is the cost of an optimal plan, and $c_{i}$ is the cost of the $i$-th plan. Clearly, the higher the SWC score is, the more optimal the resulting plans are. If all the generated plans are optimal, the SWC score reaches its maximum value 1. Last, to quantify the diversity of the generated plans, we check the number of unique correct plans out of the 64 responses for each task, and report the average number across 1000 evaluation tasks.
@@ 5.1.1 Fast Mode
Dualformer also consistently generates more diverse plans for all the considered problems. In Appendix 9, we show one Maze example and plot the unique correct plans generated by Dualformer and the baseline model. One interesting observation is that the diversity score of Dualformer, i.e., the average number of unique correct plans out of 64 responses, increases as the Maze size goes up. Intuitively, as the Maze becomes larger, there are more possible routes to reach a single goal location. This suggests that Dualformer learns the Maze structure, while the Solution-Only model is potentially memorizing the optimal plans as its diversity score is close to 1 for all the Maze sizes.
@@ 5.1.4 Ablation Study
Figure 5.1 compares the 1-Optimal-64 rate of Dualformer and Mix-$p$ models. We test 8 values of $p$: 00 (equivalent to the Complete-Trace model), $0.05,0.1,0.2,0.4,0.6,0.8$, and $1.0$ (equivalent to the Solution-Only model). In both inference modes, Dualformer beats all the Mix-$p$ models for all the five tasks we consider. In fact, Dualformer also outperforms Mix-$p$ models in the other metrics we consider. In particular, compared with all the other models that can execute in slow mode ($p<1$), it is the “fastest” one: it generates the shortest reasoning trace. Due to the space limit, we only plot the 1-Optimal-64 rate here, and refer the readers to Appendix 10 for the detailed discussion.
(a) Slow mode comparison with Mix-$p$ models.
(b) Fast mode comparison with Mix-$p$ models.
Figure 5.1: The 1-Optimal-64 rate of Dualformer and Mix-$p$ models with varying values of $p$, where $p$ is the fraction of solution-only data in the corresponding training dataset. The top and bottom panels plot the results in fast and slow mode, respectively. Dualformer outperforms all the Mix-$p$ models in both inference modes. The probabilities of different dropping strategies used for Dualformer is described in the Hyperparameter paragraph, see Section 5.1.
@@ 5.2 Application to LLM Training: Math Reasoning
We finetune two base models Mistral-7B and LLama-3-8B for two epochs using a batch size of $32$. We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $8e{-6}$ for the Mistral models, and $5e{-6}$ for the Llama-3 models. Full training details, including how we select the learning rates and other hyperparameters, are deferred to Appendix 7.2. For fine-tuning the models, we use the same prompt as in Yu et al. (2023), which is displayed in Appendix 12.2.
@@
Similar to Dualformer, we evaluate the models in both the fast and the slow mode, where the LLM is required to output the final solution directly or to solve the problem step by step. Following Yu et al. (2023), we use zero-shot prompting when evaluating the models. See Appendix 12.2 for the prompts we use in each mode. We consider the Greedy@1 metric Dubey et al. (2024); Yu et al. (2023), where for each question we generate 1 response using a temperature of 0 and verify the correctness. We also report the pass@20 metric Chen et al. (2021), where we randomly sample 20 responses using a temperature of 0.5. For reference, we also report the results of fine-tuning those models on the original MATH dataset.

### v2

@@ 1 Introduction
In psychology, the dual process theory Wason and Evans (1974) explains that thought can emerge through two distinct processes. Typically, one process is implicit (automatic and unconscious) and the other process is explicit (controlled and conscious).
The famous book Thinking, Fast and Slow Kahneman (2017) explores the concept of dual process theory extensively. It describes two different systems of thinking. System 1, corresponding to the implicit process, is fast, intuitive, and emotional. System 2, corresponding to the explicit process, is slower, more deliberative, and more logical. Kahneman applies this theory to a wide range of cognitive and behavioral activities to illustrate how these two systems influence our decision making and reasoning. As outlined in the book, System 1 can lead to biases and is prone to systematic errors due to the lack of voluntary control. As a counterbalance, System 2 is more suitable for tasks that require careful analysis, consideration, and planning. 00footnotetext: The code is available at https://github.com/facebookresearch/dualformer.
Transformers Vaswani et al. (2017), the sequence modeling tool that serves as the cornerstone of foundation models in various domains including Large Language models (LLMs) Dosovitskiy (2020); Baevski et al. (2020); Radford et al. (2021); Touvron et al. (2021); Hsu et al. (2021); Touvron et al. (2023); Dubey et al. (2024), have been widely used in many works to approach reasoning and planning problems, see e.g.,  Zhou et al. (2022); Kojima et al. (2022); Pallagani et al. (2022); Valmeekam et al. (2023a); Chen et al. (2024); Gundawar et al. (2024); Wang and Zhou (2024). Interestingly, the dual process theory generalizes to artificial intelligence, where we can categorize the reasoning modes of Transformer into fast and slow. In fast mode, a Transformer will directly output a final solution without any reasoning steps, whereas the intermediate steps of thinking, such as a search trace for finding a short path, will be generated along with the plan in slow mode. The two inference modes share similar pros and cons with the two thinking systems inherent in us. As discussed in previous works Wei et al. (2022); Valmeekam et al. (2023b); Lehnert et al. (2024); Gandhi et al. (2024); Saha et al. (2024), models operating in fast mode have a lower computational cost and allow for a quicker response, yet they fall short in accuracy and optimality compared with slow mode models.
This raises the question:
@@
We first apply our framework to train an encoder-decoder Transformer model to solve pathfinding problems, where the trace is generated by the A* search algorithm. We consider two domains: the Maze navigation and the Sokoban game as in Lehnert et al. (2024), where we use the same tokenization scheme. Interestingly, we have found that these problems are challenging for state-of-the-art LLMs like o1-preview and o1-mini, where the output path often breaks into the walls (See Section 13 for an example).
In each reasoning mode, Dualformer outperforms the established baselines, achieving stronger results in both the solved rate and optimal rate. Moreover, Dualformer significantly enhances the diversity of the generated plans by identifying a greater variety of unique paths that reach the goal. Notably, Dualformer is also efficient even when working in slow mode, generating much shorter reasoning traces than the baseline model. Next, we apply our framework to finetune LLMs for answering math questions. Following Yu et al. (2023), the training examples are taken from the MATH dataset Hendrycks et al. (2021) where answers are rewritten by a Llama-3.1-70B-Instruct model to include detailed intermediate steps. Likewise, the obtained LLMs demonstrate enhanced efficacy and efficiency.
@@ 2 Related Work
Tremendous efforts have been made to enhance the capability of Transformer-based models to plan and reason over a long horizon. Two main types of approaches have been developed. The first type leverages existing LLMs. For instance, researchers taught LLMs to call external existing symbolic solvers, such as those found in Ahn et al. (2022); Besta et al. (2024); Sel et al. (2023); He-Yueya et al. (2023); Liu et al. (2023); Silver et al. (2024). Pallagani et al. (2022, 2024) investigate the fine-tuning of LLMs to use a symbolic solver, Schick et al. ; Hao et al. (2024) fine-tune LLMs to use external tools, and Hao et al. propose to fine-tune LLMs to do reasoning and planning within a world model. Among them, several works try to integrate System-1 and System-2 thinking into LLM reasoning. Weston and Sukhbaatar (2023) proposed the System 2 Attention (S2A), a more deliberate attention mechanism aimed at improving the reasoning capabilities of large language models by reducing their susceptibility to spurious correlations in the context. Yu et al. (2024) distill System 2 output into a System 1 model, which aims to compile higher-quality outputs from System 2 techniques back into LLM generations without intermediate reasoning token sequences.
The second type aims to train Transformers from scratch to plan and reason independently Lehnert et al. (2024); Saha et al. (2024); Gandhi et al. (2024), using task-specific language. Both Lehnert et al. (2024) and Gandhi et al. (2024) teach LLMs to search by representing the search process in language, Lehnert et al. (2024) developed Searchformer that is trained to mimic the A* search process for pathfinding problems. Gandhi et al. (2024) applied their model to the Countdown game. In the concurrent work Saha et al. (2024), the authors trained two separate models and manage them by an external meta-controller. There, one model generates rapid traceless response, while another one responds slower but with trace. Similarly, Lin et al. (2023) leverages multiple models to implement slow and fast thinking using an agent style workflow. Our research is closely related to the above work, but has some key differences. Unlike Lehnert et al. (2024); Gandhi et al. (2024) which do not modify the reasoning trace in the training data, our approach involves randomizing the reasoning trace. Unlike Saha et al. (2024); Lin et al. (2023), we do not use any explicit controller nor use two networks for each mode. Instead, we integrate both the fast and the slow mode functionalities into a single model.
@@
Large language models (LLMs) have been utilized for synthetic data generation in various domains. For instance, Wei et al. (2021); Longpre et al. (2023); Taori et al. (2023) introduced a synthetic instruction dataset by sampling from diverse templates containing natural language instructions that outline specific tasks. This approach has also been applied to the visual domain Liu et al. (2024b, a); Zhu et al. (2023); Brooks et al. (2023); Peng et al. (2023). To improve the performance of LLM to answer math questions, Yu et al. (2023) developed a method to rewrite, verify, and augment the original MATH dataset Hendrycks et al. (2021) using specialized prompts. Similar methodologies have been further explored in other studies, including those by  Yuan et al. (2023); Luo et al. (2023); Lee et al. (2023); Yue et al. (2023); Tong et al. (2025).
@@ 4 Structured Trace Dropping and Randomized Training
Searchformer proposed by Lehnert et al. (2024) has proven efficacy in addressing a variety of complex decision-making tasks. However, it still suffers from two important limitations.
Firstly, the model only operates in slow mode and outputs lengthy reasoning chains, which significantly increases the inference time. While this can be reduced by bootstrapping Lehnert et al. (2024), an iterative refining technique that consists of cycles of rollouts followed by fine-tuning, such a procedure incurs significant extra demand on computational resources.
Secondly, Searchformer struggles to generate diverse solutions, where identical rollouts are frequently sampled [^1] We note that the lack of diversity is not due to the limitation of training data. For all our experiments, the search traces in the training datasets are generated by non-deterministic A* algorithm that randomly breaks cost ties.. For example, across 1000 $30\times 30$ maze problems we have tested, Searchformer’s reasoning chain contains more than 1500 tokens on average, and can only find $7.6$ unique feasible paths out of $64$ responses (see Section 5.1.2).
@@ 5.1 Navigation Tasks: Maze & Sokoban
We evaluate whether a model generates correct and optimal plans using two metrics: 1-Solved-64 and 1-Optimal-64. Namely, for each evaluation task (e.g. one maze or one Sokoban game), we randomly sample 64 responses from a trained model. Each response is parsed and evaluated regardless of the generated trace part. If any of the 64 plans is correct, i.e. is feasible and reaches the goal location, this task is labelled as success for the 1-Solved-64 metric. If any of the 64 plans is optimal, this task is labelled as success for the 1-Optimal-64 metric. We repeat such process for 1000 unseen evaluation tasks and report the average success rate. To investigate the robustness of each method, we also report metrics 3-Solved-64 and 3-Optimal-64 where a task is labelled as success if at least 3 plans are correct or optimal. Additionally, we consider the Success Weighted by Cost (SWC) Wu et al. (2019) that measures the quality of the resulting plans in terms of their costs, aggregated over individual responses. More precisely: $\text{SWC}=\tfrac{1}{64}\sum_{i=1}^{64}\mathbb{I}(\text{plan}\,i\,\text{is correct})\cdot\tfrac{c^{*}}{c_{i}}$, where $\mathbb{I}$ is the indicator function, $c^{*}$ is the cost of an optimal plan, and $c_{i}$ is the cost of the $i$-th plan. Clearly, the higher the SWC score is, the more optimal the resulting plans are. If all the generated plans are optimal, the SWC score reaches its maximum value 1. Last, to quantify the diversity of the generated plans, we check the number of unique correct plans out of the 64 responses for each task, and report the average number across 1000 evaluation tasks.
@@ 5.1.1 Fast Mode
Dualformer also consistently generates more diverse plans for all the considered problems. In Appendix 9, we show one Maze example and plot the unique correct plans generated by Dualformer and the baseline model. One interesting observation is that the diversity score of Dualformer, i.e., the average number of unique correct plans out of 64 responses, increases as the Maze size goes up. Intuitively, as the Maze becomes larger, there are more possible routes to reach a single goal location.
This suggests that Dualformer learns the Maze structure, while the Solution-Only model is potentially memorizing the optimal plans as its diversity score is close to 1 for all the Maze sizes.
@@ 5.1.3 Auto Mode: Dual Process
An interesting question to ask is whether Dualformer can automatically adjust its operation mode in response to problem difficulty. To investigate this, we generated 64 auto-mode responses for 1000 unseen mazes with varying wall densities (between 0.3 and 0.5). We then analyzed the percentage of slow-mode paths among all feasible solutions. The results are plotted in Figure 5.1. As the wall density increases, indicating a (likely) more challenging maze, the proportion of slow-mode paths also increases. This could be due to two factors: (1) the inherent need for slower thinking to solve more difficult problems, and (2) Dualformer is selecting the slow mode more frequently. Similarly, we observe that as the maze size increases, where the problem becomes harder, Dualformer consistently employs more slow thinking.
Figure 5.1: The percentage of slow mode paths (among all the feasible paths it generates) that Dualformer (auto mode) activates in. As the wall density and maze size, Dualformer engages in slow thinking more often.
5.1.4 Generalization Performance
For all the experiments we have presented so far, we test on mazes of the same size and wall density. It is intriguing to inspect the OOD generalization performance of Dualformer. In this section, we consider a Dualformer trained on mazes of size $20\times 20$, where wall density is uniform randomly sampled between $0.3$ and $0.5$. At test time, we vary the wall density from $0.1$ to $0.6$. Table 5.5 presents the slow mode results. We test 50 unseen examples for each case. As the wall density increases, we expect the maze to become more challenging and the prompt also becomes longer. Therefore, we can see for in-distribution test cases (wall density 0.3, 0.4, 0.5), Dualformer (slow mode) achieves approximately 100% optimal rate. Yet when the wall density increases to $0.6$, which is OOD, the performance drops. Interestingly, Dualformer (slow mode) does not solve a single maze for lower wall densities. Our intuition is that the prompt becomes too short for Dualformer to generalize. In addition to varying wall densities, we also check the slow mode performance of on rectangular mazes (e.g., height=20, width=10). As before, our model is trained on $20\times 20$ mazes. Table 5.6 shows the results. Surprisingly, Dualformer generalizes for the cases we consider here.
Wall Density 1-Optimal-64 1-Solved-64
0.1 (OOD) 0.0 0.0
0.2 (OOD) 0.0 0.0
0.3 97 100
0.4 100 100
0.5 100 100
0.6 (OOD) 68 72
Table 5.5: Dualformer (slow mode, trained on $20\times 20$ mazes) achieves nearly-perfect optimal rate when the wall density is in distribution, while the performance drops for OOD values.
Maze Size 1-Optimal-64 1-Solved-64
$20\times 10$ 100 100
$20\times 12$ 94 95
$20\times 14$ 100 100
$20\times 16$ 73 73
$20\times 18$ 99 100
$20\times 19$ 100 100
Table 5.6: Dualformer (slow mode, trained on $20\times 20$ mazes) generalizes to rectangular mazes.
@@ 5.1.5 Ablation Study
Figure 10.1 compares the 1-Optimal-64 rate of Dualformer and Mix-$p$ models. We test 8 values of $p$: 00 (equivalent to the Complete-Trace model), $0.05,0.1,0.2,0.4,0.6,0.8$, and $1.0$ (equivalent to the Solution-Only model). In both inference modes, Dualformer beats all the Mix-$p$ models for all the five tasks we consider. In fact, Dualformer also outperforms Mix-$p$ models in the other metrics we consider too, see Section 10. In particular, it is the “fastest” one when operating in slow mode: it generates the shortest reasoning trace.
@@ 5.2 Application to LLM Training: Math Reasoning
We finetune two base models Mistral-7B and LLama-3-8B for two epochs using a batch size of $32$. We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $8e{-6}$ for the Mistral models, and $5e{-6}$ for the Llama-3 models.
Full training details, including how we select the learning rates and other hyperparameters, are deferred to Appendix 7.2. For fine-tuning the models, we use the same prompt as in Yu et al. (2023), which is displayed in Appendix 12.2.
@@
Similar to Dualformer, we evaluate the models in both the fast and the slow mode, where the LLM is required to output the final solution directly or to solve the problem step by step. Following Yu et al. (2023), we use zero-shot prompting when evaluating the models. See Appendix 12.2 for the prompts we use in each mode. We consider the Greedy@1 metric Dubey et al. (2024); Yu et al. (2023), where for each question we generate 1 response using a temperature of 0 and verify the correctness. We also report the pass@20 metric Chen et al. (2021), where we randomly sample 20 responses using a temperature of 0.5.
For reference, we also report the results of fine-tuning those models on the original MATH dataset.